OpenStudy (anonymous):

I gave up on this question. –6p + 7 = 3(2p – 3) – 4(–10 + 4p)

OpenStudy (chaise):

Never give up :) –6p + 7 = 3(2p – 3) – 4(–10 + 4p) Expand the brackets. -6p+7 = 6p-9 +40 -16p Gather like terms. -6p+7 = -10p+31 Rearrange the equation 4p = 24 p = 6 Anything you didn't understand in this process?

OpenStudy (anonymous):

I understood somewhat, but I'll get it. Thank you. :) I'm terrible at Math.

OpenStudy (chaise):

If you still need help I don't mind explaining it for you.

OpenStudy (anonymous):

I think I've got it, but thanks for offering.